OpenStudy (anonymous):

Two angles are complementary. The measure of one angle is 21 more than twice the measure of the other angle. Find the measures of the angles.

OpenStudy (jmark):

As we know complementary angles equal 90 degrees together. Let Angle 2 = x then from given Angle 1 = 2x + 21 So, x + 2x + 21 = 90 3x + 21 = 90 3x = 69 x = 23 So Angle 2 = 23 degrees Then, you ""plug and chug"" Angle 1 = 2(23) + 21 Angle 1 = 46 + 21 Angle 1 = 67 degrees CBSE Sample Papers for Class 12 - http://cbse.edurite.com/cbse-sample-papers/cbse-sample-papers-for-class-12.html#close_iframe